# A Doppler backscattering diagnostic for the EXL-50U spherical tokamak: plasma considerations and preliminary quasioptical design

### v3

@@ Abstract
[label2]organization=Nanyang Technological University, 21 Nanyang Link, #04-01 School of Physical and Mathematical Sciences, city=Singapore, postcode=637371, country=Republic of Singapore
@@
[label3]organization=Physics and Astronomy Department, UCLA, city=Los Angeles, postcode=90098, state=California, country=USA
@@
[label4]organization=ENN Science & Technology Development Co., Ltd, city=Langfang, postcode=065001, state=Hebei, country=China
@@ 1 Introduction
The Energy iNNovation (ENN) XuanLong-50U (EXL-50U) spherical tokamak, an upgraded version of the ENN XuanLong-50 (EXL-50), was built by the ENN Energy Research Institute in China to advance vital technologies for proton-boron fusion in spherical tokamaks Liu et al. [2024], see Table 1 for its main parameters Yuejiang et al. [2025]. The EXL-50U tokamak was designed with a small aspect ratio to improve plasma confinement Liu et al. [2024]. There are two reasons for this improvement. First, a smaller aspect ratio results in greater toroidicity Kaye et al. [2021], Bingren [2003] — defined by the inverse aspect ratio Bingren [2003] — which reduces areas with bad curvature in the tokamak Kaye et al. [2021]. Secondly, the smaller aspect ratio in spherical tokamaks results in greater $E\times B$ shearing than in conventional tokamaks Kaye et al. [2021]. These two effects suppress electrostatic drift-wave instabilities at the ion and electron scales Kaye et al. [2021], thereby improving confinement. However, simulation results predicting ion- and electron-scale transport do not always agree with experimental observations Howard et al. [2016]. Cross-scale simulations that resolve both ion and electron contributions have shown that ion-scale turbulence can be indirectly amplified by electron-scale turbulence through sub-ion-scale structures Maeyama et al. [2015, 2017a], significantly reducing confinement. Hence, it is crucial to understand cross-scale turbulent interactions in spherical tokamaks such as the EXL-50U through well-resolved experimental measurements of electron- and ion-scale turbulence. Furthermore, EXL-50U uses proton-boron fusion, which requires ion temperatures to be much higher than electron temperatures Liu et al. [2024]. This regime is expected to yield novel and potentially rich turbulence physics that warrant careful experimental investigation. One such technique is Doppler backscattering (DBS), a diagnostic used to measure turbulence in many tokamaks and stellarators worldwide Hennequin et al. [2004], Happel et al. [2009], Zhou et al. [2013], Shi et al. [2016], Rhodes et al. [2016], Hu et al. [2017], Tokuzawa et al. [2018], Molina Cabrera et al. [2019], Wen et al. [2021], Tokuzawa et al. [2021], Carralero et al. [2021], Yashin et al. [2021], Rhodes et al. [2022], Shi et al. [2023], Macwan et al. [2024], Conway et al. [2025b]. DBS is able to measure flows Pratt et al. [2022] and density fluctuations of intermediate length scales, usually 1 $\lesssim$ $k_{\perp}\rho_{s}$ $\lesssim$ 10, where $\rho_{s}$ is the deuterium ion sound gyroradius, which depends on electron temperature. Furthermore, it can measure turbulent fluctuations in the plasma core, which is challenging.
@@
Here, $l$ is the arc length along the central ray, $F_{m}$ is the mismatch attenuation, $F_{i}$ is the non-mismatch part of the instrumentation function, and $\tilde{n}_{e}$ is the amplitude of the electron density fluctuations as a function of the backscattered wavenumber, which is given by the Bragg condition at each point along the central ray. This is a simplified version of the backscattered spectral density from Hall-Chen et al. [2022b], which is sufficient for this paper. The mismatch attenuation is given by
@@
This results in mismatch attenuation, which is a decrease in the backscattered power because the turbulence wavevector, which is perpendicular to the magnetic field vector, is not aligned with the probe beam’s wavevector. A larger mismatch attenuation means that there will be no signal, which is a bigger problem at larger $K$ values Hall-Chen et al. [2022b]. This is important to note in this work because we want to measure turbulent wavenumbers $k_{\perp}$} that are high enough to be considered electron scale. Furthermore, a decrease in backscattered power could either mean a decrease in turbulent fluctuations or an increase in the mismatch angle between the normal to the magnetic field and the wavevector of the probe beam, and these causes are experimentally indistinguishable just from looking at the signal alone.
In cases where the poloidal magnetic field is significant compared to the toroidal magnetic field, such as in tokamaks with high magnetic pitch angles, the incident beam’s wavevector may not be perpendicular to the magnetic field. Hence, understanding mismatch attenuation is important in this work. The EXL-50U tokamak has a high magnetic pitch angle (Fig. 3), which means that the mismatch can be large if the toroidal launch angles are not carefully chosen. However, it is possible to minimise the mismatch by using toroidal steering to vary the toroidal launch angle Damba et al. [2022], Hall-Chen et al. [2022c]. Hence, we will need to determine the optimal toroidal launch angle that yields zero mismatch for a given beam frequency. For $\Delta\theta_{m}$, the beam width and curvature values are needed to calculate it, and these depend on the design of the DBS system used to propagate a beam of a particular beam width and curvature.
@@
The rest of the paper is structured as follows: we present the plasma scenarios used in this study in Section 2 along with the estimated range of probe-beam frequencies required. We then use SCOTTY, in ray-tracing mode, to calculate the cutoff locations and scattered wavenumbers as a function of poloidal launch angle, see Section 3. We then design the quasioptical system to achieve this range of frequencies and poloidal launch angles, subject to physical constraints; refer to Section 4. In Section 5, we use beam-tracing to determine the mismatch attenuation and therefore the toroidal steering required. We also estimate the spatial resolution of high-$k$ measurements. Finally, we suggest potential further improvements in Section 6 and conclude in Section 7.
@@ 3.2.2 H-mode (B) cutoff locations
Although the plasma core is accessed over a wide poloidal flux range, the probed locations are sparse. For X-mode beams above 50 GHz and O-mode beams above 42 GHz, cutoff locations for different beam frequencies at a given poloidal launch angle are more dispersed than at lower frequencies, where they cluster closely, see Fig. 7(b) and Fig. 8(b) for X-mode, and Fig. 7(e) and Fig. 9(b) for O-mode. When $f_{r}>50~\textrm{GHz}$ and $f_{pe}>42~\textrm{GHz}$, $f_{r}$ and $f_{pe}$ change slowly with position, so a slight change in frequency causes a considerable change in cutoff location (see Fig. 4(b)). This leads to a large spatial separation between the cutoff points, limiting plasma core coverage. To probe density fluctuations between cutoff points in the core, measurements can be sampled at smaller frequency intervals, increasing spatial and poloidal flux coverage. As in H-mode (A), this can be done using a tunable frequency channel.
@@ 4 Designing the quasioptics
Having determined the frequencies and poloidal launch angles in the previous section, we now proceed to design the quasioptics for the proposed EXL-50U DBS. This design must satisfy certain constraints, such as having sufficient clearance from the probe beam to propagate. As a starting point, we consider an ex-vessel system consisting of a scalar horn antenna, a high-density polyethylene (HDPE) biconvex lens to focus the probe beam and thus reduce its width, and a steering mirror, see Fig. 10. This design is similar to DBS diagnostics in the literature Hillesheim et al. [2015].
Figure 10: Overview of the DBS used to send microwaves into the plasma. The antenna emits microwaves at various frequencies, which are then focused by a HDPE biconvex lens before being reflected into the plasma by a flat mirror with two-axis steering that sets the initial launch angles of the beam.
@@ 4.1 Requirements
    The beam width at the mirror must be less than 1/3 the diameter of the mirror. Hence, we set the mirror radius to be 1.5 times the beam width at the mirror.
@@
    The beam width at the lens must be less than 1/3 the diameter of the lens.
@@ 4.2 Proposed quasioptical system design
    Type of horn and its far-field divergence, given by the full width at half maximum (FWHM) angle.
@@
With the beam properties defined at the entry, the lens’s focal length and the lens-mirror distance constrain the remaining two parameters. To estimate these two parameters, we propagate 40 GHz and 60 GHz beams with the aforementioned beam properties from the plasma entry, back out to the vacuum vessel, through the lens and then to the horn. The focal length and lens-mirror distance are varied to determine the optimal location for the horn, fixing the third parameter: horn-lens distance.
The beam widths of the 40 GHz and 60 GHz beams at the horn are then used to determine the horn properties needed. The 40 GHz and 60 GHz beams correspond to the maximum and minimum beam widths at the horn location, respectively. Hence, we take the mean of the two widths $w_{m}$ to determine the horn aperture radius $a$ and the FWHM angle $\theta_{FWHM}$ Goldsmith and others [1998]:
$a=\frac{w_{m}}{0.644},$ (24)
and
$\theta_{FWHM}=\sqrt{2\ln{2}}\tan^{-1}\left(\frac{c}{\pi w_{m}f_{mid}}\right).$ (25)
Here, $c$ is the speed of light and $f_{mid}$ is the mid-band frequency taken as 50 GHz. This determines the last parameter.
Now that we have the four parameters, we propagate the beam from the horn, through the lens and into the plasma. Next, to check whether requirement 5 can be satisfied, the lens properties must be calculated. As mentioned in requirement 5, the lens radius can be calculated by multiplying the maximum beam width at the lens by 1.5. The radius of curvature of the lens is determined using the thin lens lensmaker’s formula, given by,
@@
where $f$ is the focal length of the lens, $n$ is the refractive index of HDPE, and $R_{1}$ and $R_{2}$ are the radii of curvature of the lens. Note that we have assumed a frequency-independent refractive index of 1.52 for HDPE. Since we are using a biconvex lens, $R_{1}=-R_{2}$, so we can determine the radius of curvature of the lens given by,
@@
Having found the design parameters of our quasioptical components (horn, lens, mirror, and distances between them), we give them in Table 5.
@@
FWHM beamwidth $8.23^{\circ}$
@@
HDPE refractive index 1.52
@@ 4.3 Port window design
Figure 13: (a) Diagram showing the quantities used to calculate the beam widths at the port window and the port window diameter. (b) The port window is large enough to account for not only the beam widths but also the separation between the beams.
@@ 5.1.1 H-mode (A)
For the X-mode beams, the optimal toroidal launch angle does not vary significantly with frequency, so $\varphi_{t}=5^{\circ}$ can be used to avoid severe mismatch, see Fig. 14(b) and Fig. 14(e). The toroidal launch angle for zero mismatch remains nearly constant because the beam reaches cutoff at about the same location in the pedestal for any frequency from 40 GHz to 60 GHz, see Fig. 4.
However, the optimal toroidal launch angle can generally vary with frequency, as seen in the O-mode case, where it has a noticeable variation in the optimal toroidal launch angle from 54 GHz onwards, see Fig. 15(b) and Fig. 15(e). This is because when an O-mode beam is propagated with frequencies above 54 GHz, it penetrates the pedestal and reaches cutoff at a very different location, see Fig. 4. Hence, for a given toroidal launch angle, some frequencies will be matched while others will be mismatched, and the measurable frequency range for each toroidal angle varies and typically does not span the entire 20 GHz range. Therefore, toroidal steering with a tunable frequency channel should be used to prevent severe mismatch.
@@ 5.1.2 H-mode (B)
The optimal toroidal launch angle for X-mode beams stays nearly constant from 40 GHz to 52 GHz, then decreases from about $8^{\circ}$ to $1^{\circ}$ between 52 GHz and 58 GHz, see Fig. 14(a) and Fig. 14(d). This is because X-mode beams from 40 GHz to 52 GHz reaches cutoff at roughly the same location in the pedestal region. Beams above these frequencies reach cutoff at a different location, which explains the large change in toroidal angle that yields zero mismatch. This matches Fig. 4, where X-mode beams from 40 GHz to 50 GHz reach cutoff at a similar radial coordinate in the pedestal. In comparison, beams with frequencies above 50 GHz can penetrate the pedestal.
The trend is similar for O-mode. The optimal toroidal angle also stays nearly constant from 40 GHz to 48 GHz, but decreases from about $8^{\circ}$ to $1^{\circ}$ between 48 GHz and 54 GHz, see Fig. 15(a) and Fig. 15(d). Low-frequency O-mode beams also reaches cutoff at a similar radial coordinate, whereas higher frequency beams can reach cutoff at different locations, consistent with Fig. 4. As in H-mode (A), there is no single compromise angle that yields significant backscattering from both the edge and core channels. Toroidal steering with a tunable frequency channel should be used to prevent beams from becoming severely mismatched.

### v4

@@ Abstract
[label4]organization=Physics and Astronomy Department, UCLA, city=Los Angeles, postcode=90098, state=California, country=USA
[…]
[label5]organization=ENN Science & Technology Development Co., Ltd, city=Langfang, postcode=065001, state=Hebei, country=China
@@ 1 Introduction
The Energy iNNovation (ENN) XuanLong-50U (EXL-50U) spherical tokamak, an upgraded version of the ENN XuanLong-50 (EXL-50), was built by the ENN Energy Research Institute in China to advance vital technologies for proton-boron fusion in spherical tokamaks Liu et al. [2024], see Table 1 for its main parameters Yuejiang et al. [2025]. The EXL-50U tokamak was designed with a small aspect ratio to improve plasma confinement Liu et al. [2024]. There are two reasons for this improvement. First, a smaller aspect ratio results in greater toroidicity Kaye et al. [2021], Bingren [2003] — defined by the inverse aspect ratio Bingren [2003] — which reduces areas with bad curvature in the tokamak Kaye et al. [2021]. Secondly, the smaller aspect ratio in spherical tokamaks results in greater $E\times B$ shearing than in conventional tokamaks Kaye et al. [2021]. These two effects suppress electrostatic drift-wave instabilities at the ion and electron scales Kaye et al. [2021], thereby improving confinement. However, simulation results predicting ion- and electron-scale transport do not always agree with experimental observations Howard et al. [2016]. Cross-scale simulations that resolve both ion and electron contributions have shown that ion-scale turbulence can be indirectly amplified by electron-scale turbulence through sub-ion-scale structures Maeyama et al. [2015, 2017a], significantly reducing confinement. In this study, we aim to understand cross-scale turbulent interactions in spherical tokamaks such as the EXL-50U through well-resolved experimental measurements of ion-scale turbulence and, in certain cases, the lower end of electron scale turbulence. Hence, this DBS potentially enables the investigation of cross-scale turbulent interactions. Furthermore, EXL-50U uses proton-boron fusion, which requires ion temperatures to be much higher than electron temperatures Liu et al. [2024]. This regime is expected to yield novel and potentially rich turbulence physics that warrant careful experimental investigation. One such technique is Doppler backscattering (DBS), a diagnostic used to measure turbulence in many tokamaks and stellarators worldwide Hennequin et al. [2004], Happel et al. [2009], Zhou et al. [2013], Shi et al. [2016], Rhodes et al. [2016], Hu et al. [2017], Tokuzawa et al. [2018], Molina Cabrera et al. [2019], Wen et al. [2021], Tokuzawa et al. [2021], Carralero et al. [2021], Yashin et al. [2021], Rhodes et al. [2022], Shi et al. [2023], Macwan et al. [2024], Conway et al. [2025b]. DBS is able to measure flows Pratt et al. [2022] and density fluctuations of intermediate length scales, usually 1 $\lesssim$ $k_{\perp}\rho_{s}$ $\lesssim$ 10, where $\rho_{s}$ is the deuterium ion sound gyroradius, which depends on electron temperature. Furthermore, it can measure turbulent fluctuations in the plasma core, which is challenging.
@@
Here, $l$ is the arc length along the central ray, $F_{m}$ is the mismatch attenuation, and $F_{i}$ is the non-mismatch part of the instrumentation function. $\tilde{n}_{e}$ is the amplitude of the electron density fluctuations as a function of the backscattered wavenumber, which is given by the Bragg condition at each point along the central ray. This is a simplified version of the backscattered spectral density from Hall-Chen et al. [2022b], which is sufficient for this paper. The mismatch attenuation is given by
@@
Hence, when the probe beam’s wavevector is not perpendicular to the magnetic field vector, the backscattered power is reduced due to mismatch attenuation. A larger mismatch attenuation means that there will be no signal, which is a bigger problem at larger $K$ values Hall-Chen et al. [2022b]. This is important to note in this work because we want to measure turbulent wavenumbers $k_{\perp}$} that are high enough to be considered electron scale. Furthermore, a decrease in backscattered power could either mean a decrease in turbulent fluctuations or an increase in the mismatch angle between the normal to the magnetic field and the wavevector of the probe beam, and these causes are experimentally indistinguishable just from looking at the signal alone.
In cases where the poloidal magnetic field is significant compared to the toroidal magnetic field, such as in tokamaks with high magnetic pitch angles, the incident beam’s wavevector may not be perpendicular to the magnetic field. Hence, understanding mismatch attenuation is important in this work. The EXL-50U tokamak has a high magnetic pitch angle (Fig. 3), which means that the mismatch can be large if the toroidal launch angles are not carefully chosen. However, it is possible to use toroidal steering to minimise mismatch for a given beam frequency and poloidal launch angle Damba et al. [2022], Hall-Chen et al. [2022c]. Having steering in both the poloidal and toroidal directions (2D steering) also allows us to potentially make magnetic pitch angle measurements Yeoh et al. [2026]. For $\Delta\theta_{m}$, the beam width and curvature values are needed to calculate it, and these depend on the design of the DBS system used to propagate a beam of a particular beam width and curvature.
@@
The rest of the paper is structured as follows: we present the plasma scenarios used in this study in Section 2 along with the estimated range of probe-beam frequencies required. We then use SCOTTY, in ray-tracing mode, to calculate the cutoff locations and scattered wavenumbers as a function of poloidal launch angle, see Section 3. We then design the quasioptical system to achieve this range of frequencies and poloidal launch angles, subject to physical constraints; refer to Section 4. In Section 5, we use beam-tracing to determine the mismatch attenuation and therefore the toroidal steering required, and also estimate the spatial resolution of high-$k$ measurements. We finally conclude in Section 6.
@@ 3.2.2 H-mode (B) cutoff locations
Although the plasma core is accessed over a wide poloidal flux range, the probed locations are sparse. For X-mode beams above 50 GHz and O-mode beams above 42 GHz, cutoff locations for different beam frequencies at a given poloidal launch angle are more dispersed than at lower frequencies, see Fig. 7(b) and Fig. 8(b) for X-mode, and Fig. 7(e) and Fig. 9(b) for O-mode. When $f_{r}>50~\textrm{GHz}$ and $f_{pe}>42~\textrm{GHz}$, $f_{r}$ and $f_{pe}$ change slowly with position, so a slight change in frequency causes a considerable change in cutoff location (see Fig. 4(b)). This leads to a large spatial separation between the cutoff points, limiting plasma core coverage. To probe density fluctuations between cutoff points in the core, measurements can be sampled at smaller frequency intervals, increasing spatial and poloidal flux coverage. As in H-mode (A), this can be done using a tunable frequency channel.
@@ 4 Designing the quasioptics
Having determined the frequencies and poloidal launch angles in the previous section, we now proceed to design the quasioptics for the proposed EXL-50U DBS. This design must satisfy certain constraints, such as having sufficient clearance from the probe beam to propagate. As a starting point, we consider an ex-vessel system consisting of a scalar horn antenna, an ultra high molecular weight polyethylene (UHMWPE) biconvex lens to focus the probe beam and thus reduce its width, and a steering mirror, see Fig. 10. This design is similar to DBS diagnostics in the literature Hillesheim et al. [2015].
Figure 10: Overview of the DBS used to send microwaves into the plasma. The antenna emits microwaves at various frequencies, which are then focused by a UHMWPE biconvex lens before being reflected into the plasma by a flat mirror with two-axis steering that sets the initial launch angles of the beam.
@@ 4.1 Requirements
    The beam width at the lens must be less than 1/3 the diameter of the lens.
@@
    The projected beam width at the mirror must be less than 1/3 the diameter of the mirror. Hence, we set the mirror radius to be 1.5 times the projected beam width at the mirror.
@@ 4.2 Proposed quasioptical system design
    Aperture radius of the horn.
@@
With the beam properties defined at the entry, the lens’s focal length and the lens-mirror distance constrain the remaining two parameters. To estimate these two parameters, we propagate 40 GHz and 60 GHz beams with the aforementioned beam properties from the plasma entry, back out to the vacuum vessel, through the lens and then to the horn. The focal length and lens-mirror distance are varied to determine the optimal location for the horn, fixing the third parameter: horn-lens distance. The beam widths of the 40 GHz and 60 GHz beams at the horn are then used to determine the horn aperture radius $a$ needed. The 40 GHz and 60 GHz beams correspond to the maximum and minimum beam widths at the horn location, respectively. Hence, we take the mean of the two widths $w_{m}$ to determine $a$ Goldsmith and others [1998]. We obtain
$a=\frac{w_{m}}{0.644}.$ (24)
Now that we have the four parameters, we propagate the beam from the horn, through the lens and into the plasma. Next, to check whether requirement 4 can be satisfied, the lens properties must be calculated. As mentioned in requirement 4, the lens radius can be calculated by multiplying the maximum beam width at the lens by 1.5. The radius of curvature of the lens is determined using the thin lens lensmaker’s formula, given by,
@@
where $f$ is the focal length of the lens, $n$ is the refractive index of UHMWPE, and $R_{1}$ and $R_{2}$ are the radii of curvature of the lens. Note that we have assumed a frequency-independent refractive index of 1.575 for UHMWPE Xie et al. [2020]. Since we are using a biconvex lens, $R_{1}=-R_{2}$, so we can determine the radius of curvature of the lens given by,
@@
Finally, we determine the radius of the mirror. The projected width on the mirror $w_{projected}$ is given by,
$w_{projected}=\frac{w_{mirror}}{\cos{\Phi}}.$ (27)
Here, $w_{mirror}$ is the 40 GHz beam width at the mirror, and $\Phi$ is the angle of incidence, defined as the angle between the central ray and the normal of the mirror. $\Phi$ corresponds to the incidence angle when the beam is launched at $\varphi_{p}=-35.5^{\circ}$. To minimise the projected width, $\Phi$ has to be minimised. Hence, the system is oriented such that the antenna launches the beam vertically downward onto the mirror, see Fig. 12(a). This configuration yields $\Phi=27.3^{\circ}$.
Figure 12: (a) Diagram showing the orientation of the quasioptical system, and the quantities used to calculate the beam widths at the port window and the port window diameter. (b) The port window is large enough to account for not only the beam widths but also the separation between the beams.
The mirror radius can be calculated by multiplying $w_{projected}$ by 1.5.
Additionally, the far-field divergence of the horn, given by the full width at half maximum (FWHM) angle $\theta_{FWHM}$, can be computed. This is given by Goldsmith and others [1998],
$\theta_{FWHM}=\sqrt{2\ln{2}}\tan^{-1}\left(\frac{c}{\pi w_{m}f_{beam}}\right).$ (28)
Here, $c$ is the speed of light and $f_{beam}$ is the beam frequency. Note that $\theta_{FWHM}$ varies with frequency. Having found the design parameters of our quasioptical components (horn, lens, mirror, and distances between them), we give them in Table 5.
@@
FWHM beamwidth $6.88^{\circ}$$10.27^{\circ}$
@@
UHMWPE refractive index 1.575
@@ 5.1.1 H-mode (A)
For the X-mode beams, the optimal toroidal launch angle does not vary significantly with frequency, so $\varphi_{t}=5^{\circ}$ can be used to avoid severe mismatch, see Fig. 14(a) and Fig. 14(d). The toroidal launch angle for zero mismatch remains nearly constant because the beam reaches cutoff at about the same location in the pedestal for any frequency from 40 GHz to 60 GHz, see Fig. 4.
However, the optimal toroidal launch angle can generally vary with frequency, as seen in the O-mode case, where it has a noticeable variation in the optimal toroidal launch angle from 54 GHz onwards, see Fig. 15(a) and Fig. 15(d). This is because when an O-mode beam is propagated with frequencies above 54 GHz, it penetrates the pedestal and reaches cutoff at a very different location, see Fig. 4. Hence, for a given toroidal launch angle, some frequencies will be matched while others will be mismatched, and the measurable frequency range for each toroidal angle varies and typically does not span the entire 20 GHz range. Therefore, toroidal steering with a tunable frequency channel should be used to prevent severe mismatch.
@@ 5.1.2 H-mode (B)
The optimal toroidal launch angle for X-mode beams stays nearly constant from 40 GHz to 52 GHz, then decreases from about $8^{\circ}$ to $1^{\circ}$ between 52 GHz and 58 GHz, see Fig. 14(b) and Fig. 14(e). This is because X-mode beams from 40 GHz to 52 GHz reaches cutoff at roughly the same location in the pedestal region. Beams above these frequencies reach cutoff at a different location, which explains the large change in toroidal angle that yields zero mismatch. This matches Fig. 4, where X-mode beams from 40 GHz to 50 GHz reach cutoff at a similar radial coordinate in the pedestal. In comparison, beams with frequencies above 50 GHz can penetrate the pedestal.
The trend is similar for O-mode. The optimal toroidal angle also stays nearly constant from 40 GHz to 48 GHz, but decreases from about $8^{\circ}$ to $1^{\circ}$ between 48 GHz and 54 GHz, see Fig. 15(b) and Fig. 15(e). Low-frequency O-mode beams also reaches cutoff at a similar radial coordinate, whereas higher frequency beams can reach cutoff at different locations, consistent with Fig. 4. As in H-mode (A), there is no single compromise angle that yields significant backscattering from both the edge and core channels. Toroidal steering with a tunable frequency channel should be used to prevent beams from becoming severely mismatched.